Determine the area under the standard normal curve that lies to the left of ​(a) z equals 0.73 comma ​(b) z equals negative 0.54​, ​(c) z equals negative 0.14​, and​ (d) z equals 1.65.
Answer #1

We have to calculate area under standard normal curve that is left of given z-scores.

That is we have to calculate P( Z < z) = ?

a)

We have to calculate P( Z < 0.73 ) = ?

From Z table,

P( Z < 0.73) = 0.7673

b)

P( Z < -0.54) = 1 - P( Z < 0.54)

= 1 - 0.7054 (from Z table)

= 0.2946

c)

P( Z < -0.14) = 1 - P( Z < 0.14)

= 1 - 0.5557 (From Z table)

= 0.4443

d)

P( Z < 1.65) = 0.9505 (From Z table)